# Multi-Modal Artificial Intelligence of Embryo Grading and Pregnancy Prediction in Assisted Reproductive Technology: A Review

### v1

@@ IV-D Public dataset
Since medical image data usually involve patient privacy information and legal compliance requirements, public datasets may violate relevant regulations or cannot pass ethical review, so most ART-related datasets have not been opened to the public. Considering the importance of open data to promote the development of medical artificial intelligence and computer-assisted therapy technology, this paper summarizes the currently publicly available representative datasets as listed in TABLE IV.
In addition to these public datasets, some open models such as STORK are also widely used for embryo evaluation. STORK (https://github.com/ih-lab/STORK) is an AI model from Stanford University, trained on the Inception image recognition architecture. The model can identify the morphological characteristics of the embryo and output the transfer probability or classification results. Other commercial software scoring systems, such as KIDScore and iDAScore [111], are also widely used. They are fully AI-driven scoring systems developed by Vitrolife, which are trained on time-lapse imaging sequences without manual labeling, outputting a continuous score (e.g., 1 to 10) for prioritization of embryos. Meanwhile, ERICA (Embryo Ranking Intelligent Classification Algorithm) [112] is also an embryo evaluation model based on deep learning, aiming to automatically score and rank static embryos in ART, which can be accessed on https://embryoranking.com/.
V Challenge and future work
@@
V-A Multi-modal fusion
In the process of ART, data types show a high degree of heterogeneity, including static image, time-lapse video, structured clinical data, and genomic information. Multi-modal data fusion is considered to be a key strategy to improve model prediction performance and achieve comprehensive embryo quality assessment. However, there are still several bottlenecks in the actual modeling process. First, there are differences in dimensionality and semantics between data, which makes it difficult to achieve effective fusion. So, the redundancy and even conflict in multi-modal data may reduce the accuracy of the model. Second, the complex structure of the fusion model leads to large computing resource consumption in the training process, and it is difficult to deploy in resource-constrained scenarios.
For the multi-modal fusion task, future research can focus on the design of lightweight fusion methods. For fusion strategies, in order to be able to effectively fuse the features among multi-modalities, common operation-based fusion methods such as AddFusion, ConcatFusion, and TensorFusion [113] can be used [114]. Furthermore, GNN, with their message passing mechanism based on graph structure, can establish effective connections between different modal features in multi-modal data, thereby mining high-order semantic relationships among cross-modal nodes and achieving information fusion [115]. At the same time, the Transformer architecture can effectively model the long-term dependence within a single modality through the self-attention mechanism, and its cross-attention mechanism further promotes the information interaction and alignment between different modalities [116]. The model can maintain the semantic consistency and complementarity between modalities in the fusion process. Compared with other fusion methods, decoupled fusion method decomposes multi-modal features into shared features and unique features, avoids redundant interference between modalities, and improves the model’s understanding of each modality [78][117][118][119].
The problem of lightweight model can be solved by using lightweight framework. For the image modality, MobileNet [120] and ShuffleNet [121] series networks can be used, and for the tabular modality, a simple MLP can be used. It can also be solved by knowledge distillation [122], using a large model (Teacher) to guide a small model (Student) to learn an intermediate representation or output distribution.
V-B Data resources
The data related to assisted reproduction have the characteristics of high sensitivity to human privacy, strict ethical approval requirements, and difficulty in sample collection, resulting in the general small datasets. And most of the experimental results reported in this paper are based on private datasets, which makes the results incomparable and hinders the technical development of AI in this field. In particular, there are still significant deficiencies in the construction of multi-modal datasets involving embryo morphological images and supporting clinical records. This problem directly limits the use of AI models in real-world scenarios.
In order to solve the problem of scarcity of public high-quality datasets in the field of ART, the most direct way is to collect clinical data manually. However, this method faces many institutional challenges in practice and is difficult to be widely promoted. In view of the difficulty of model training caused by the limited amount of data, a variety of effective alternative paths have been proposed in current research. Generative models such as GAN [123], Diffusion [124] can be used to synthesize datasets, thus increasing the diversity of datasets [105]. In addition, Few-shot Learning also provides an idea to deal with the small sample size problem. By introducing metric learning or meta-learning mechanisms, few-shot learning enables the model to achieve better performance with only a small number of labeled samples [125].
V-C Model generalization
@@
In order to improve the cross-domain adaptation ability of the model, the future should focus on the development of transfer learning [126] and domain adaptation [127] technology to realize the dynamic adjustment of model parameters or feature space. At the same time, as a collaborative training framework with local data, federated learning [128] provides a feasible path for breaking data islands and realizing cross-institution model sharing.
V-D Ethics and law
AI’s decision-making participation in ART, especially when it involves sensitive tasks such as embryo scoring, genetic screening, and pregnancy success prediction, raises a series of ethical and legal issues. For example, the scoring of embryos may be seen as judging the ”potential value of life”. In addition, AI intervention in the field of reproduction may also have a subtle impact on society’s conception of fertility and family structure ethics.
Therefore, it is urgent to establish a systematic ethical evaluation framework and technical supervision mechanism to clarify the responsibility of AI in assisted reproduction. And relevant practitioners should develop norms covering data privacy protection, informed consent, model interpretability, and transparency of use.
VI Conclusion
@@
Acknowledgment
This work is supported in part by the Guangdong Provincial Natural Science Foundation (2023A1515011431), the Guangzhou Science and Technology Planning Project (202201010092), the National Natural Science Foundation of China (72074105).

### v2

@@ 6 Public Dataset
Since medical image data usually involve patient privacy information and legal compliance requirements, public datasets may violate relevant regulations or cannot pass ethical review, so most ART-related datasets have not been opened to the public. Considering the importance of open data to promote the development of medical artificial intelligence and computer-assisted therapy technology, this paper summarizes the currently publicly available representative datasets as listed in TABLE 5.
Although some datasets in the field of ART have been published, their distribution and quality are significantly limited. Specifically, most of the existing public datasets focus on the embryo grading task, which has been maturely applied and widely implemented in clinical practice, and its research value tends to be saturated [143]. In contrast, the dataset of embryo segmentation task is not only released earlier, but also limited in sample size, which is difficult to meet the requirements of current deep learning models for data quantity and quality. In addition, for the critical clinical task of multi-modal pregnancy prediction, only a small dataset is available [155], which is far from enough to support the training and validation of the model. Therefore, there is an urgent need to construct high-quality and large-scale pregnancy prediction datasets to promote the scientific research and clinical translation application of related algorithms.
7 Challenge and Future Work
@@
7.1 Multi-modal fusion
As discussed in Section 4.3, the task of embryo grading has reached a state of maturity, achieving successful clinical translation [143]. The sign of embryo grading maturity lies in the fact that its performance has surpassed that of professional embryologists [42][43][47]. Consequently, the focus of contemporary research is progressively shifting towards the more formidable challenge of predicting pregnancy outcomes. The inherent biological complexity of pregnancy, governed by a multitude of interacting factors, cannot be fully captured by models reliant on uni-modal data, such as static embryo images. This limitation constitutes a fundamental performance bottleneck. Therefore, the integration of multi-modal information—encompassing static images, time-lapse videos, and structured clinical and laboratory data—has emerged as a pivotal strategy and a frontier direction for enhancing the predictive accuracy of these models.
Despite its considerable potential, the practical implementation of multi-modal fusion strategies is impeded by two core bottlenecks. The first challenge stems from the heterogeneity and semantic gap inherent in multi-modal data [156]. Static images, time-lapse videos, and structured tabular data originate from distinct feature spaces and exhibit significant disparities in their intrinsic semantic representations and information densities. A naive approach, such as simple feature concatenation or early fusion, often results in a “forced alignment” that fails to achieve effective cross-modal synergy. Instead, this method risks introducing noise and degrading model performance due to semantic conflicts or information redundancy between modalities, a phenomenon known in the literature as negative transfer [157].
The second bottleneck pertains to the architectural complexity and substantial computational overhead of multi-modal models [158]. These models typically necessitate parallel feature extraction networks and sophisticated fusion modules, resulting in an architecture that is significantly larger and more complex than their uni-modal counterparts. This complexity translates into a formidable demand for computational resources, including high-performance GPUs, large memory capacity, and protracted training times. This high computational cost not only constrains the efficiency of model development and iterative optimization but also poses a significant barrier to the practical deployment of such models in clinically resource-constrained environments.
To address the aforementioned challenges, future research in pregnancy prediction for ART should prioritize two key directions: the innovation of multi-modal fusion strategies and the architectural design of lightweight models.
7.1.1 Multi-modal Fusion Strategies
Traditional fusion methods, such as simple feature concatenation or element-wise operations, are computationally efficient but fail to capture the complex, non-linear interactions inherent in heterogeneous ART data [159][160]. To overcome this limitation, more sophisticated paradigms are required. GNNs offer a powerful framework by modeling multi-modal entities as nodes in a graph, enabling the explicit mining of high-order semantic relationships through message-passing mechanisms [161]. Similarly, the Transformer architecture, with its self-attention and cross-attention mechanisms, excels at capturing long-range dependencies within individual modalities (e.g., time-lapse videos) and facilitating precise cross-modal information alignment [162]. Furthermore, decoupled fusion presents a promising approach to mitigate negative transfer by decomposing multi-modal features into modality-invariant and modality-specific subspaces, thereby preserving complementary information while minimizing interference [108][163][164][165].
7.1.2 Model Lightweight
While these advanced fusion strategies enhance predictive performance, they often introduce substantial computational complexity, hindering deployment in resource-constrained clinical settings. Therefore, model lightweight is imperative. A primary approach involves employing lightweight backbone networks [166][167], such as MobileNet for image data or simplified MLPs for tabular data, which utilize efficient architectural designs like depth-wise separable convolutions to reduce the parameter count and computational load. A more sophisticated technique is knowledge distillation, where a compact “student” model is trained to get the ability of a larger, high-performing “teacher” model [168][169]. The resulting highly accurate and efficient models are essential for enabling real-time AI applications in clinical ART.
7.2 Data resources
As discussed in Section 6, data resources in the field of ART exhibit significant scarcity characteristics due to their inherent sensitivity to human privacy, ethical review protocols, and complex sample collection processes. This data dilemma precipitates two interrelated challenges that critically impede the robust iteration and objective evaluation of AI models in ART domain.
Firstly, the lack of reproducibility and benchmarking undermines scientific progress. At present, the vast majority of published research results are based on the private data sets of each institution. Due to the differences in data distribution, collection standards and processing procedures, it is difficult to compare and evaluate the experimental results between different studies, which greatly weakens the universality of research conclusions. Especially, the field of ART still lacks a large-scale, high-quality, standardized public multi-modal dataset, and this ”data gap” has become a key bottleneck for AI to move from algorithm research to clinical application in this field.
Secondly, the adaptation of existing models on private small datasets is insufficient. When faced with limited scale and expensive labeling of clinical data, directly transferring models pre-trained on general large-scale datasets, or relying on simple data augmentation techniques, often yields little benefit. These methods are difficult to fundamentally solve the problem of overfitting caused by insufficient data, so it is urgent to develop more sophisticated algorithm strategies designed for small sample scenarios.
To address the above challenges, academia and industry have explored a number of parallel and complementary technical paths. At the data level, although it is the most ideal solution to manually collect and construct large-scale public datasets through multi-center cooperation, it is often difficult to be widely promoted in the short term due to legal barriers to data sharing, interest coordination between institutions, and long ethical approval processes. Thus, leveraging generative models to ”create” data becomes a highly potential alternative. For example, generative adversarial networks [170] and emerging diffusion models [171] can learn the intrinsic distribution of real data and generate highly realistic and diverse synthetic embryo images, thus effectively expanding the size and diversity of the training set and alleviating the problem of data scarcity under the premise of protecting patient privacy. There have been relevant works [153][172] applying generation technology to the field of ART.
At the algorithmic level, few-shot learning provides a fundamental paradigm for solving the small sample size problem. It enables the model to ”learn how to learn” by introducing mechanisms such as metric learning or meta-learning [173]. The combination of these advanced algorithm strategies provides practical technical support for breaking through the data barriers in the field of ART and promoting the clinical implementation of AI models.
7.3 Model generalization
@@
In order to improve the cross-domain adaptation ability of the model, the future should focus on the development of transfer learning [174] and domain adaptation [175] technology to realize the dynamic adjustment of model parameters or feature space. At the same time, as a collaborative training framework with local data, federated learning [176] provides a feasible path for breaking data islands and realizing cross-institution model sharing. For example, Wang et al. [177] proposed a federated learning framework combined with an expectation maximization algorithm for image assessment of embryonic cleavage quality, which enhanced prediction accuracy while protecting privacy.
7.4 Ethics and law
AI’s decision-making participation in ART, especially when it involves sensitive tasks such as embryo scoring, genetic screening, and pregnancy success prediction, raises a series of ethical and legal issues. For example, the scoring of embryos may be seen as judging the ”potential value of life”. In addition, AI intervention in the field of reproduction may also have a subtle impact on society’s conception of fertility and family structure ethics. Among them, the most prominent challenge is the question of responsibility [178]. With the continuous digitalization of the healthcare system and the wide application of AI algorithms in clinical decision support, more and more non-traditional medical agents, including technology developers, data providers, and algorithm operators, are gradually involved in the medical service chain, thus reshaping the subject structure of medical responsibility [179]. However, when an AI system leads to misdiagnosis, treatment delay, or other adverse clinical consequences due to algorithmic flaws or data bias, the existing legal framework faces a significant dilemma in defining who is responsible: should the developer, medical institution, clinical user, or multiple parties share the responsibility? Consensus has not yet emerged on this issue [180].
To complicate matters, most advanced AI models (especially deep learning systems) have a ”black box” nature that lacks transparency and traceability in their decision-making processes [181]. Although Explainable Artificial Intelligence (XAI) methods have made certain progress in recent years, their explanations often stay at the surface correlation, and it is difficult to reveal the causal mechanism inside the model. This inherent opacity not only weakens clinicians’ ability to understand and judge the output results of AI, but also shakes their trust in the system, which affects the effectiveness and safety of human-computer collaborative decision-making [182].
In addition, the behavioral deviation in human-computer interaction further aggravates the ethical and legal risks. Research shows that experienced clinicians tend to be cautious or even skeptical of AI recommendations and tend to retain independent judgment; Junior providers, on the other hand, may rely too much on system outputs and suffer from ”automation bias,” which is to blindly trust technology over clinical intuition and comprehensive assessment [183]. This cognitive difference not only affects the quality of diagnosis and treatment, but also introduces uncertainty in the interaction between human and algorithm in the responsibility chain, which makes the accident imputation more complex.
Therefore, it is urgent to establish a systematic ethical evaluation framework and technical supervision mechanism to clarify the responsibility of AI in assisted reproduction. And relevant practitioners should develop norms covering data privacy protection, informed consent, model interpretability, and transparency of use. Although nascent initiatives for ethical guidelines and policy-making have begun to emerge, the governance frameworks for artificial intelligence technologies significantly lag behind their rapid advancement in the medical domain [184]. Concurrently, there is a prevailing lack of comprehensive awareness within the medical community regarding the complex ethical challenges posed by these emerging AI technologies [185]. Consequently, the risk assessment of AI-powered medical tools should not be confined solely to algorithmic accuracy. Instead, it necessitates a more holistic and context-specific evaluation. For instance, the U.S. Food and Drug Administration (FDA) typically mandates that approved AI medical devices be maintained in a “locked” state, wherein their algorithms are prohibited from being modified post-market [186]. This regulatory paradigm serves to reinforce product safety and stability.
8 Conclusion
@@
[…]
List of Abbreviations
| ACC | Accuracy |
| AFC | Antral Follicle Count |
| AI | Artificial Intelligence |
| AMH | Anti-Mullerian Hormone |
| ART | Assisted Reproductive Technology |
| AUC | Area Under receiver operating characteristic Curve |
| BC | BlastCoele |
| BG | Background |
| BMI | Body Mass Index |
| CC | Cell Counting |
| CNN | Convolutional Neural Network |
| EG | Embryo Grading |
| EMT | EndoMetrial Thickness |
| GNN | Graph Neural Network |
| Grad-CAM | Gradient-weighted Class Activation Mapping |
| hCG | Human Chorionic Gonadotropin |
| ICM | Inner Cell Mass |
| ICSI | IntraCytoplasmic Sperm Injection |
| IVF-ET | In Vitro Fertilization-Embryo Transfer |
| LIME | Local Interpretable Modelagnostic Explanations |
| PGT | Preimplantation Genetic Testing |
| PP | Pregnancy Prediction |
| RNN | Recurrent Neural Network |
| SHAP | SHapley Additive exPlanations |
| SSL | Self-Supervise Learning |
| TE | TrophEctoderm |
| TLM | Time-Lapse Microscopy |
| t-SNE | t-distributed Stochastic Neighbor Embedding |
| ZP | Zona Pellucida |
Appendix A Background
A.1 The first generation——IVF-ET
Step 5: The selected embryo is transferred into the mother’s uterus.
Day 1: The zygote starts the first cleavage, forming a 2-cell embryo.
Day 2: The embryo continues cleavage at the 2-4 cell stage.
Day 3: The embryo continue cleavage at the 6-8 cell stage.
A.2 The second generation——ICSI
A.3 The third generation——PGT
Author Biographies
Xueqiang Ouyang is an Assistant Engineer at the Information Department, The People’s Hospital of Baoan Shenzhen, China. He received his Bachelor’s and Master’s degrees from the South China University of Technology in 2022 and 2025. His main research is Artificial Intelligence in Medicine.
Jia Wei is an Associate Professor at the School of Computer Science and Engineering, South China University of Technology, China. He received his Bachelor’s and Master’s degrees in Computer Science from the Harbin Institute of Technology in 2003 and 2006, and his Ph.D. from the South China University of Technology in 2009. His main research interests include Machine Learning, Deep Learning, Artificial Intelligence in Medicine.